OpenStudy (anonymous):

Brownsville is planting trees along both sides of its streets. the trees will be planted 50 ft apart. the first tree on the block must stand 20 ft from the corner. if the blocks in Brownsville are 390 ft long how many trees will be planted along one block? A.) 16 trees B.) 7 trees C.) 8 trees D.) 14 trees

OpenStudy (anonymous):

if Im not mistaken the ans is D... the block is 390 feet long the 1st tree is 20 ft away from the corner so 390-20=370 feet then all trees 50 ft apart so every 50 ft u plant a tree so 370/50= 7.4 so thats 7 trees and its 2 sides of the street 7*2=14... hopefully I did that right ;)